# NuHepMC: A standardized event record format for neutrino event generators Steven Gardiner1, Joshua Isaacson1 and Luke Pickering2

### v2

@@ Acknowledgements
Appendix A Target Scaling Conventions
It is useful to illustrate the relationship between the different cross-section target scaling conventions discussed in G.C.4 with an example. Let and respectively denote the total cross section for a given process to occur with a and atom. Then the cross section for an interaction with the carbon part and the hydrogen part of a molecule under the different target scaling conventions are given below:
"PerTargetAtom":
Carbon interaction:
Hydrogen interaction:
"PerTargetMolecule":
Carbon interaction:
Hydrogen interaction:
"PerTargetNucleon":
Carbon interaction:
Hydrogen interaction:
"PerTargetMolecularNucleon":
Carbon interaction:
Hydrogen interaction:
It should be clear from this example why we recommend against implementations using "PerTargetMolecularNucleon", important information about the relative cross section for the nuclear constituents is lost when averaging over the molecule—the cross-section information reported for an interaction with the hydrogen part of a hydrocarbon will implicitly contain carbon nuclear effects. However, the scheme is discussed here because where cross-section information is written out at the event level in existing neutrino event generators, it is often used. It is usually possible to reconstruct the cross-section for constituents by brute force from a sample of events.
Appendix B Estimation via Monte Carlo sampling
In this section we provide two example methods of estimating the flux-averaged total cross-section when it is not known a priori. The first method is suitable for GENIE-like codes that calculate energy-dependent total cross sections prior to event generation itself. The second is suitable for ACHILLES and similar codes that do not use precalculated cross sections in this way.
B.1 GENIE-like approach
@@
It follows from these definitions that
@@
Note that there is no universally accepted expression for the standard error on a weighted arithmetic mean such as the one computed in Eq. B.3 (before taking the reciprocal of the expression). The choice used in Eq. B.4 is based on the third expression for “” recommended in Ref. [30] based on bootstrapping studies (see also references therein). The quantity proposed as an estimator for the square of the standard error on a weighted mean
@@
The expression in Eq. B.6 may be simplified to read
The result in Eq. B.4 is obtained immediately with the substitutions and .
While the expressions in Eqs. B.3 and B.4 may be useful for estimating from a sample of generated events, they require access to the entire sample and are thus unsuitable for implementing the running estimate described in E.C.4. However, an adaptation of West’s algorithm [31] provides a solution for this application. Let
@@
B.2 ACHILLES-like approach
@@
The results of the above Monte-Carlo calculation would produce a set of weighted events that can be used as is, or can be unweighted through the following procedure. First, the maximum values for each neutrino species and nuclei can be estimated using Monte-Carlo methods (). Second, a neutrino species and nuclei is selected according to the probability
where is given by the sum of the maximum weight over all neutrino types and nuclei. Once a neutrino type and nucleus is selected, a set of initial and final state momenta are generated and the integrand for that particular neutrino type and nucleus is calculated. The event can then be unweighted by performing an accept-reject step using the ratio of weight of the event to the maximum weight calculated for the given neutrino type and nucleus. In other words, the weight of an event would be given by
@@
Appendix C Example Event Graphs
Figs. 2, 3, and 4 show some example event graphs for neutrino event generators that natively implement, or are convertible-to, the NuHepMC format. Although the details of each generator’s implementation of this standard differ, the constraints that are imposed enable consistent usage for the most common tasks without any knowledge of each generator’s implementation details.
Appendix D Extracting Bibliography Information
If the NuHepMC file signals that convention G.C.6 has been used, then it is possible to run the file through an external tool called HEPReference [35]111The code can be found at: https://github.com/NuDevTools/HEPReference to produce a BibTeX file and a short text blurb showing how to cite all the papers used to produce the events. An example use case can be found in the repository.

### v3

@@ Acknowledgements
Appendix A Estimation via Monte Carlo sampling
In this section we provide two example methods of estimating the flux-averaged total cross-section when it is not known a priori. The first method is suitable for GENIE-like codes that calculate energy-dependent total cross sections prior to event generation itself. The second is suitable for ACHILLES-like codes that do not use pre-calculated total cross sections in this way.
A.1 GENIE-like approach
@@
It follows from the definitions in Sec. 3 that
@@
Note that there is no universally accepted expression for the standard error on a weighted arithmetic mean such as the one computed in Eq. A.3 (before taking the reciprocal of the expression). The choice used in Eq. A.4 is based on the third expression for “” recommended in Ref. [32] based on bootstrapping studies (see also references therein). The quantity proposed as an estimator for the square of the standard error on a weighted mean
@@
The expression in Eq. A.6 may be simplified to read
The result in Eq. A.4 is obtained immediately with the substitutions and .
While the expressions in Eqs. A.3 and A.4 may be useful for estimating from a sample of generated events, they require access to the entire sample and are thus unsuitable for implementing the running estimate described in E.C.4. However, an adaptation of West’s algorithm [33] provides a solution for this application. Let
@@
A.2 ACHILLES-like approach
@@
The results of the above Monte-Carlo calculation would produce a set of weighted events that can be used as is, or can be unweighted through the following procedure. First, the maximum values for each neutrino species and nucleus can be estimated using Monte-Carlo methods (). Second, a neutrino species and nucleus is selected according to the probability
where is given by the sum of the maximum weight over all neutrino types and nuclei. Once a neutrino type and nucleus is selected, a set of initial and final state momenta is generated and the integrand for that particular neutrino type and nucleus is calculated. The event can then be unweighted by performing an accept-reject step using the ratio of weight of the event to the maximum weight calculated for the given neutrino type and nucleus. In other words, the weight of an event would be given by
@@
Appendix B Example Event Graphs
Figs. 2, 3, 4, 5 and 6 show some example event graphs for neutrino event generators that natively implement, or are convertible-to, the NuHepMC format. Although the details of each generator’s implementation of this standard differ, the constraints that are imposed enable consistent usage for the most common tasks without any knowledge of each generator’s implementation details.
Appendix C Extracting Bibliography Information
If the NuHepMC file signals that convention G.C.3 has been used, then it is possible to run the file through an external tool called HEPReference [37]999The code can be found at: https://github.com/NuDevTools/HEPReference to produce a BibTeX file and a short text blurb showing how to cite all the papers used to produce the events. An example use case can be found in the repository.